Countries Mali and South Africa have their interest rates to be 16% and 12 %, respectively. If their currencies trade according to 50 CFA francs buy one rand in the spot market, what will their future spot rate be in the aforementioned context?
Answer #1

Firstly,

Spot rate means its the current market price for which the asset (commodity or security or a currency) is bought or sold for immediate delivery.

Formula for future spot rate = spot rate * (1+foreign interest rate)/(1+domestic interest rate)

= 50 * (1+16%) / (1+12%)

= 50 * (1.16)/(1.12)

= 50* 1.0357

= 51.785